Topic: Integration



Find the antiderivative of the following. Show your comple solution.



1.) Integral (6x⁵-4x²+2) dx


2.) Integral (12x²+2x) dx


3.) Integral 8^3/2 dx


4.) Integral 1/x¹² dx


5.) Integral ³√x dx

1
Expert's answer
2022-05-03T16:41:34-0400

1)

"\int(6x^5-4x^3+2) dx=6(\dfrac{x^6}{6})-4(\dfrac{x^4}{4})+2x+C"

"=x^6-x^4+2x+C"

2)

"\int(12x^2+2x) dx=12(\dfrac{x^3}{3})+2(\dfrac{x^2}{2})+C"

"=4x^3+x^2+C"

3)

"\int(8^{3/2}) dx=8^{3/2}\int dx"

"=8^{3\/2}x+C"

4)

"\int(\dfrac{1}{x^{12}}) dx=\int x^{-12}dx=\dfrac{x^{-11}}{-11}+C"

"=-\dfrac{1}{11x^{11}}+C"

5)

"\int \sqrt[3]{x}dx=\int x^{1/3}dx=\dfrac{x^{4/3}}{4/3}+C"

"=\dfrac{3x\sqrt[3]{x}}{4}+C"